### v1

# LEREL: Lipschitz Continuity-Constrained Emotion Recognition Ensemble Learning For Electroencephalography

Abstract
Accurate and efficient perception of emotional states in oneself and others is crucial, as emotion-related disorders are associated with severe psychosocial impairments. While electroencephalography (EEG) offers a powerful tool for emotion detection, current EEG-based emotion recognition (EER) methods face key limitations: insufficient model stability, limited accuracy in processing high-dimensional nonlinear EEG signals, and poor robustness against intra-subject variability and signal noise. To address these challenges, we propose LEREL (Lipschitz continuity-constrained Emotion Recognition Ensemble Learning), a novel framework that significantly enhances both the accuracy and robustness of emotion recognition performance. The LEREL framework employs Lipschitz continuity constraints to enhance model stability and generalization in EEG emotion recognition, reducing signal variability and noise susceptibility while maintaining strong performance on small-sample datasets. The ensemble learning strategy reduces single-model bias and variance through multi-classifier decision fusion, further optimizing overall performance. Experimental results on three public benchmark datasets (EAV, FACED and SEED) demonstrate LEREL’s effectiveness, achieving average recognition accuracies of 76.43%, 83.00% and 89.22%, respectively.
Keywords Electroencephalography (EEG)  $\cdot$ Lipschitz Continuity  $\cdot$ Ensemble Learning  $\cdot$ Emotion Recognition
1 Introduction
Emotion recognition – the capacity to assess emotional states in oneself and others – is fundamental to social functioning. This critical ability is impaired in conditions such as autism spectrum disorder, depression and schizophrenia, where deficits are strongly associated with adverse psychosocial outcomes including social withdrawal, self-harm behaviors, and suicidal ideation. The perceptual efficiency and accuracy of emotion recognition therefore represent vital targets for both clinical intervention and computational modeling approaches. Electroencephalography (EEG)-based emotion recognition (EER) technology shows great potential for mental health applications, enabling objective assessment and diagnosis of emotional disorders to facilitate early intervention and precision treatment strategies [1] [2]. Emotions are complex, dynamic psychological states shaped by interpersonal and neurobiological factors. The study of emotions poses significant research challenges, particularly in emotion recognition. While no universal consensus exists on emotion categorization, individual differences in emotional expression, especially among individuals with autism and depression, are critically important for both theoretical and clinical applications [3]. This underscores the need for objective emotion assessment tools capable of precisely capturing internal emotional states to inform clinical diagnosis and guide personalized therapeutic interventions.
Recent advances in EER have focused on developing more accurate [4] [5] and robust [6] models. Promising approaches include multimodal data fusion of EEG with other physiological signals [7] [8] [9] [10] [11] [12] [13] and the construction of brain network features to capture complex signal patterns [14]. Despite these innovations, EER still faces significant challenges, including poor signal-to-noise ration, susceptibility to artifacts, high intra-individual variability [15] [16] [17] [18] [19], and limitations in model generalizability.
Lipschitz continuous constraints provide a rigorous mathematical framework to enhance model stability by restricting gradient variations, ensuring outputs remain robust to minor perturbations. For example, Jia et al. demonstrated improved node-level adversarial robustness in graph neural networks through Lipschitz regularization, significantly enhancing stability with noisy data [20]. Their subsequent work applied similar constraints in a fairness regularization method to control output variations and ensure individual fairness [21]. These constraints also strengthen generalization by regularizing model behavior across diverse data distributions and individual variations. This makes Lipschitz continuous constraints particularly valuable for addressing EER challenges, where signal variability and noise sensitivity are persistent issues [22] [23] [21].
Figure 1: Overall framework of the LEREL
In this study, to address these challenges in EER, we propose a Lipschitz continuity-constrained Emotion Recognition Ensemble Learning (LEREL) framework with the following key contributions:
  • The integration of Lipschitz continuity constraints enhances model stability and generalization during training, effectively addressing EEG signal variability while improving robustness and recognition accuracy.
  • The LEREL framework employs an ensemble learning strategy that harnesses the synergistic strengths of multiple classifiers, significantly enhancing both the precision and resilience of emotion recognition.
  • The LEREL framework demonstrates outstanding performance in emotion recognition across multiple benchmark datasets, confirming its effectiveness and generalizability.
2 Method
2.1 Problem Definition
Generally speaking, the EEG signal can be expressed as
$\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ (1)
where $\mathbf{x}_{T}\in\mathbb{R}^{C}$ represents the signal vector at time point $t$, $C$ is the number of EEG channels, and $T$ is the length of the time series. Further, we can define the emotional category set as following:
$\mathcal{Y}=\{y_{1},y_{2},\dots,y_{K}\}$ (2)
where $K$ is the total number of emotional categories. Thus, we can define the EER problem as a task that aims to discern an individual’s emotional state at a particular moment or over a temporal span based on EEG data. The corresponding mathematic expression can formulated as the emotional category label $\hat{y}$ prediction process based on the given $\mathbf{X}$, following the maximization of the prediction probability as:
$\hat{y}=\arg\max_{y_{k}}P(Y=y_{k}\mid\mathbf{X})$ (3)
2.2 Lipschitz Continuity
Lipschitz continuity is a property that describes the rate of change of a function and is strictly defined in metric space. Incorporating Lipschitz continuity into models can effectively enhance their generalization and robustness, making it particularly valuable for EEG signal processing where noise and variability are commonly challenging [24][25][26][27].
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two metric spaces, where $d_{X}$ and $d_{Y}$ are the distance functions on $X$ and $Y$, respectively. A function $f:X\to Y$ is said to be Lipschitz continuous if there exists a non-negative constant $L$ such that for any two points $x_{1},x_{2}\in X$, the following inequality holds:
$d_{Y}(f(x_{1}),f(x_{2}))\leq L\cdot d_{X}(x_{1},x_{2})$ (4)
This inequality indicates the extent to which the function $f$ “amplifies” or “reduces” the distance between any two points in $X$. For real-valued functions $f:\mathbb{R}\to\mathbb{R}$, Lipschitz continuity can be expressed as:
$|f(x_{1})-f(x_{2})|\leq L|x_{1}-x_{2}|$ (5)
This inequality implies that the absolute value of the slope of the line connecting any two points on the graph of the function will not exceed $L$. In other words, the “steepness” of the function is bounded by $L$. For example, the linear function $f(x)=ax+b$ is Lipschitz continuous with a Lipschitz constant of $|a|$. If $f$ is differentiable on the interval $[a,b]$ and its derivative $f^{\prime}(x)$ is bounded, i.e., there exists a constant $M$ such that $|f^{\prime}(x)|\leq M$ for all $x\in[a,b]$, then $f$ is Lipschitz continuous on $[a,b]$ with a Lipschitz constant $L\leq M$.
2.3 Lipschitz Gradient-Constrained Band Extraction Component (LGCBE)
The Lipschitz Gradient-Constrained Band Extraction (LGCBE) component is an advanced neural network module designed to process complex signal data. It integrates band decomposition, attention mechanisms, and Lipschitz constraints to enhance signal quality and extract key features. By capturing higher-order relationships and dynamically adjusting modality weights, it effectively handles complex interactions and contextual nuances, improving the accuracy and robustness of signal processing tasks.
Figure 2: The implementation of LGCBE
The input signal $\mathbf{X}$ has dimensions with $[B,C,T]$, where $B$ is the batch size, $C$ is the number of channels, and $T$ is the sequence length. First, the input signal is transformed from the time domain to the frequency domain using the Fast Fourier Transform (FFT):
$\mathbf{F}=\text{FFT}(\mathbf{X},\text{dim}=-1)$ (6)
The resulting frequency-domain signal $\mathbf{F}$ has dimensions with $[B,C,F]$, where $F$ is the dimension of the frequency domain. For each frequency band $b$ (defined according to standard EEG frequency bands such as delta, theta, alpha, beta, and gamma), the band mask is computed and band features are extracted as:
$\mathbf{M}_{b}=\mathbb{I}_{\text{low}_{b}\leq f\leq\text{high}_{b}}$ (7)
$\mathbf{F}_{b}=\mathbf{F}\odot\mathbf{M}_{b}$ (8)
where $\mathbb{I}$ is the indicator function and $\odot$ denotes element-wise multiplication. The energy of each frequency band is calculated as:
$\mathbf{P}_{b}=\sum_{f}|\mathbf{F}_{b}|^{2}$ (9)
The resulting band energy $\mathbf{P}_{b}$ has dimensions $[B,C]$. The channel attention weights are computed as:
$\mathbf{W}_{c}=\sigma\left(\text{MLP}_{\text{channel}}\left(\frac{1}{|\text{bands}|}\sum_{b\in\text{bands}}\mathbf{P}_{b}\right)\right)$ (10)
where $\sigma$ is the Sigmoid activation function and $\text{MLP}_{\text{channel}}$ is a network comprising two fully connected layers. The spectral attention scores are computed as:
$\mathbf{S}_{s}=\text{Softmax}\left(\text{MLP}_{\text{spectral}}\left(\frac{1}{C}\sum_{c=1}^{C}\mathbf{P}_{b}\right)\right)$ (11)
where $\text{MLP}_{\text{spectral}}$ is a network comprising two fully connected layers.
To satisfy the Lipschitz continuity constraints, the spectral attention scores are normalized:
$\mathbf{W}_{s}=\frac{\mathbf{S}_{s}}{\|\mathbf{S}_{s}\|+\epsilon}\times L$ (12)
where $\epsilon=10^{-6}$ is a numerical stability constant and $L$ is the Lipschitz constant. Then, the enhanced signal is reconstructed by combining the band features through weighted summation:
$\mathbf{X}_{\text{combined}}=\sum_{b\in\text{bands}}\mathbf{F}_{b}\odot\mathbf{W}_{c}\odot\mathbf{W}_{s,b}$ (13)
where $\mathbf{W}_{s,b}$ is the expanded form of the spectral weights. The signal is transformed back to the time domain using the Inverse Fast Fourier Transform (IFFT):
$\mathbf{y}=\text{IFFT}(\mathbf{X}_{\text{combined}},n=T,\text{dim}=-1)$ (14)
and normalization is applied:
$\mathbf{y}=\text{LayerNorm}(\mathbf{y})$ (15)
To enhance model stability, the spectral normalization is applied to the weights of the linear layers:
$\mathbf{W}=\frac{\mathbf{W}}{\|\mathbf{W}\|_{2}}\times L$ (16)
where $\|\mathbf{W}\|_{2}$ is the spectral norm of the weight matrix.
Through the above steps, the LGCBE component effectively enhances signals and extracts key features while ensuring model stability and generalization through Lipschitz continuity constraints.
2.4 Lipschitz Gradient-Constrained Attention Component (LGCA)
In recent years, the attention mechanism has made significant progress in the field of deep learning, especially in handling sequential data and graph-structured data. The Transformer architecture has achieved efficient parallel computation and long-range dependency modeling through self-attention mechanisms [28]. However, the standard dot-product self-attention mechanism tends to exhibit overconfidence when dealing with unbounded input domains, meaning that the model provides high-confidence predictions even when it is uncertain [29]. Moreover, deep Graph Attention Networks (GAT) are prone to gradient explosion when increasing the number of layers, which limits the model’s performance in tasks involving long-range dependencies [30].
To address these issues, researchers have proposed several improvements. For example, Ye et al. introduced Lipschitz regularization to mitigate the overconfidence of Transformer models [29]. Dasoulas et al. proposed a Lipschitz normalization method for self-attention layers, significantly improving the performance of deep Graph Neural Networks [30]. Lipschitz continuity regularization enhances neural networks’ robustness and generalization ability while stabilizing the training process [31, 32]. Inspired by these studies, we designed a Lipschitz Gradient-Constrained Attention component (LGCA) to enhance the model’s stability, robustness, and generalization ability by restricting its Lipschitz constant [33].
Given an input tensor $\mathbf{x}\in\mathbb{R}^{B\times T\times D}$, where $B$ is the batch size, $T$ is the sequence length, and $D$ is the embedding dimension, the conventional attention mechanism is implemented through the following steps:
The input tensor $\mathbf{x}$ is mapped to the Query, Key, and Value spaces through linear transformations:
$\mathbf{Q}=\mathbf{W}^{Q}\mathbf{x}+\mathbf{b}^{Q},\quad\mathbf{K}=\mathbf{W}^{K}\mathbf{x}+\mathbf{b}^{K},\quad\mathbf{V}=\mathbf{W}^{V}\mathbf{x}+\mathbf{b}^{V}$ (17)
where $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}$ are the weight matrices, and $\mathbf{b}^{Q},\mathbf{b}^{K},\mathbf{b}^{V}$ are the bias vectors.
The dot product is computed between the query and key, and scaled by the factor $\sqrt{d_{h}}$ to prevent the dot product results from becoming too large, which can be expressed as:
$\mathbf{S}=\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{h}}}$ (18)
where $\mathbf{S}$ denote the scaled dot-product scores and $d_{h}$ is the dimension of each attention head.
The attention scores $\mathbf{S}$ are normalized to a probability distribution using the Softmax function as follow:
$\mathbf{P}=\text{softmax}(\mathbf{S})$ (19)
where $\mathbf{P}$ denote the resulting attention weights. Then, the dropout operation is applied to the attention weights $\mathbf{P}$ to prevent overfitting:
$\mathbf{P}^{\prime}=\text{dropout}(\mathbf{P})$ (20)
Next, the value tensor $\mathbf{V}$ is combined with the attention weights $\mathbf{P}\textquoteright$ ’ through weighted summation, and the multi-head results are merged into the final output through the output projection layer. Let $\mathbf{Y}$ denote the final output:
$\mathbf{Y}=\mathbf{W}^{O}(\mathbf{P}^{\prime}\mathbf{V})+\mathbf{b}^{O}$ (21)
where $\mathbf{W}^{O}$ and $\mathbf{b}^{O}$ are the weight matrix and bias vector of the output projection layer, respectively.
Figure 3: The implementation of LGCA
Figure 4: The implementation of LGCN
2.5 Lipschitz Gradient-Constrained Normalization Component (LGCN)
Given an input tensor $\mathbf{X}\in\mathbb{R}^{B\times D}$, where $B$ denotes the batch size and $D$ denotes the feature dimension, the Lipschitz constant $L$ is used to control the gradient changes after normalization, and $\epsilon$ is a small constant to prevent division by zero.
We first perform the initial normalization of the input tensor by subtracting the mean and dividing by the standard deviation:
$\mathbf{X}_{\text{norm}}=\frac{\mathbf{X}-\mathbf{\mu}}{\mathbf{\sigma}+\epsilon}$ (22)
where $\mathbf{\mu}$ and $\mathbf{\sigma}$ are the mean and standard deviation of $\mathbf{X}$, respectively.
To introduce the Lipschitz continuity constraint, we calculate the gradient norm of the normalized tensor:
$g=\|\mathbf{X}_{\text{norm}}\|$ (23)
Then, we compute the Lipschitz constraint factor to ensure that the gradient norm does not exceed the Lipschitz constant:
$\lambda=\max\left(1,\frac{g}{L}\right)$ (24)
Here, $\lambda$ is a scalar used to adjust $\mathbf{X}_{\text{norm}}$ so that its gradient norm does not exceed $L$.
Finally, we apply the Lipschitz constraint by adjusting the normalized tensor:
$\mathbf{X}_{\text{lip}}=\frac{\mathbf{X}_{\text{norm}}}{\lambda+\epsilon}$ (25)
This step ensures that the gradient changes of $\mathbf{X}_{\text{lip}}$ are bounded by the Lipschitz constant $L$, which improves training stability and enhances the model’s generalization capability when processing complex EEG signals.
3 Experiments
3.1 Datasets
To comprehensively evaluate our LEREL, we utilized three representative EEG-based emotion recognition datasets.
The EAV dataset [34] provides a multimodal resource for affective computing research, combining synchronized EEG, audio, and video recordings from 42 participants engaged in naturalistic conversational interactions. This dataset captures five distinct emotional states—neutral, anger, happiness, sadness, and calmness—and comprises 8,400 annotated records, establishing itself as a pioneering benchmark in emotion recognition studies.
The FACED dataset [35] advances EEG-based emotion recognition research by leveraging 32-channel electrophysiological recordings from 123 participants exposed to 28 emotionally evocative video stimuli. This comprehensive resource spans nine discrete emotion categories and is further enriched with subjective self-assessment labels for emotional states, alongside multidimensional affective parameters such as arousal and valence.
The SEED dataset [36] comprises high-resolution 62-channel EEG recordings and synchronized eye movement data acquired from 15 participants exposed to emotionally evocative film clips. These stimuli were meticulously designed to elicit three distinct affective states: positive, negative, and neutral. The experimental protocol involved three sessions per subject, each containing 15 trials.
3.2 Experimental Results
In this study, the original data were resampled to 100 Hz to reduce data volume and improve efficiency while retaining key features according to the Nyquist theorem. The resampled data were segmented every 500 samples to extract local features and reduce complexity. The stratified sampling was used to construct training and testing sets, ensuring consistent sample distribution and improving model stability and generalization. The Adam [37] optimizer was employed during model training stage.
On the EAV dataset, we compared LEREL with the reported results in [34], and AMERL [9]. Since AMERL is a multimodal-based emotion recognition model that used active signals (i.e., signals collected when subjects were engaged in conversational stimuli) in its experimental validation, and our method only used EEG signals for learning, we focused on comparing the EEG-based recognition performance of AMERL. LEREL achieved a recognition accuracy of 76.43% and an F1-score of 76.22% on the EAV dataset, significantly outperforming existing methods and demonstrating superior EER performance. The corresponding results are presented in Table 1 and Figure 7.
Table 1: The results of EAV
EAV AMERL-EEG LEREL (ours) EAV AMERL-EEG LEREL (ours)
Subject ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%) Subject ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%)
1 67.50 - 59.17 - 75.83 75.80 23 63.30 - 50.00 - 76.67 75.67
2 80.00 - 64.17 - 89.17 89.10 24 75.00 - 76.67 - 88.33 88.34
3 52.50 - 54.17 - 80.00 79.97 25 49.20 - 50.00 - 77.50 78.03
4 77.50 - 66.67 - 81.67 81.34 26 58.30 - 49.17 - 73.33 72.92
5 52.50 - 40.00 - 65.00 65.11 27 70.00 - 65.00 - 86.67 86.51
6 47.50 - 48.33 - 79.17 79.35 28 63.30 - 67.50 - 87.50 87.39
7 59.20 - 59.17 - 87.50 87.29 29 51.70 - 42.50 - 62.50 61.96
8 56.70 - 54.17 - 83.33 83.33 30 59.20 - 50.00 - 69.17 68.81
9 44.20 - 45.00 - 78.33 77.81 31 48.30 - 44.17 - 74.17 73.95
10 66.70 - 47.50 - 70.83 70.68 32 60.00 - 55.83 - 71.67 71.36
11 46.70 - 44.17 - 82.50 82.68 33 73.30 - 62.50 - 84.17 84.05
12 65.80 - 45.00 - 79.17 78.13 34 54.20 - 46.67 - 67.50 66.58
13 64.20 - 55.00 - 71.67 72.17 35 30.80 - 28.33 - 44.17 43.93
14 47.50 - $3$ 66.67 65.97 36 58.30 - 60.83 - 79.17 79.37
15 59.20 - 51.67 - 70.83 70.67 37 54.20 - 55.83 - 75.00 75.04
16 55.00 - 42.50 - 71.67 71.81 38 64.20 - 45.83 - 77.50 77.34
17 78.30 - 67.50 - 84.17 84.14 39 59.20 - 50.83 - 80.83 80.46
18 69.20 - 64.17 - 84.17 83.90 40 40.00 - 40.00 - 59.17 58.75
19 55.00 - 51.67 - 75.00 74.95 41 26.70 - 43.33 - 64.17 63.15
20 77.50 - 73.33 - 87.50 87.29 42 72.50 - 65.00 - 84.17 83.89
21 68.30 - 60.00 - 85.83 86.07 Average 59.51 - 53.51 - 76.43 76.22
22 76.70 - 70.83 - 76.67 76.28 - - - - - - -
On the FACED dataset, we evaluated LEREL against the original benchmark results in [35]. LEREL achieved 83.00% recognition accuracy and 72.70% F1-score, substantially exceeding baseline performance metrics and demonstrating enhanced robustness in processing heterogeneous EEG signals. The corresponding results are presented in Table 2 and Figure 7.
Table 2: The results of FACED
Subject ACC(%) Subject ACC(%) Subject ACC(%) Subject ACC(%) Subject ACC(%)
1 92.06 26 84.13 51 72.22 76 82.54 101 86.51
2 87.30 27 76.19 52 86.51 77 88.89 102 85.71
3 81.75 28 85.71 53 93.65 78 84.13 103 80.16
4 88.89 29 98.41 54 81.75 79 89.68 104 79.37
5 66.67 30 88.89 55 95.24 80 87.30 105 75.40
6 73.81 31 86.51 56 95.24 81 87.30 106 91.27
7 87.30 32 80.16 57 90.48 82 80.95 107 78.57
8 79.37 33 64.29 58 80.16 83 93.65 108 84.13
9 89.68 34 72.22 59 95.24 84 78.57 109 84.92
10 85.71 35 93.65 60 90.48 85 83.33 110 83.33
11 72.22 36 86.51 61 65.87 86 76.19 111 77.78
12 65.87 37 73.81 62 90.48 87 90.48 112 86.51
13 84.92 38 85.71 63 82.54 88 84.13 113 85.71
14 80.95 39 90.48 64 92.86 89 89.68 114 78.57
15 76.98 40 71.43 65 95.24 90 76.98 115 84.92
16 72.22 41 86.51 66 69.84 91 67.46 116 96.03
17 71.43 42 66.67 67 77.78 92 84.13 117 87.30
18 79.37 43 84.92 68 81.75 93 80.16 118 81.75
19 74.60 44 80.16 69 90.48 94 81.75 119 81.75
20 89.68 45 69.84 70 87.30 95 86.51 120 82.54
21 86.51 46 87.30 71 93.65 96 91.27 121 95.24
22 85.71 47 81.75 72 82.54 97 89.68 122 80.16
23 80.95 48 81.75 73 80.16 98 100.00 123 75.40
24 48.41 49 95.24 74 87.30 99 95.24 Average 83.00
25 93.65 50 61.11 75 95.24 100 79.37 - -
Figure 5: Confusion Matrix of EAV
Figure 6: Confusion Matrix of FACED
Figure 7: Confusion Matrix of SEED
On the SEED dataset, we further validated the performance of our LEREL. The experimental results showed that LEREL achieved an accuracy of 89.22% and an F1-score of 89.06%, which substantiated the efficacy of LEREL across diverse emotion recognition paradigms. The corresponding results are presented in Table 3 and Figure 7.
Table 3: The results of SEED
Subject ACC(%) F1(%) Subject ACC(%) F1(%) Subject ACC(%) F1(%)
1-1 91.14 92.23 6-1 92.61 89.74 11-1 90.18 87.16
1-2 87.75 85.93 6-2 94.58 94.97 11-2 89.93 89.07
1-3 78.30 80.60 6-3 90.24 92.46 11-3 92.55 92.45
2-1 80.10 78.80 7-1 83.36 86.50 12-1 88.45 91.81
2-2 80.68 83.29 7-2 89.59 89.96 12-2 89.42 88.80
2-3 85.98 86.98 7-3 90.45 88.74 12-3 91.95 91.53
3-1 90.03 89.66 8-1 87.74 88.42 13-1 88.02 86.67
3-2 90.00 87.73 8-2 89.49 91.60 13-2 89.70 91.32
3-3 89.48 85.86 8-3 94.29 91.98 13-3 90.70 91.27
4-1 93.70 88.67 9-1 92.79 94.67 14-1 93.42 93.74
4-2 82.62 80.57 9-2 93.41 90.19 14-2 83.15 83.27
4-3 78.22 79.90 9-3 88.32 85.32 14-3 87.92 85.66
5-1 92.53 92.88 10-1 87.23 87.39 15-1 96.33 97.87
5-2 90.34 92.70 10-2 87.55 85.28 15-2 92.65 94.90
5-3 94.28 94.02 10-3 86.08 87.52 15-3 97.55 97.55
Average 89.22 89.06 - - - - - -
In summary, we conducted rigorous experiments across multiple representative datasets to evaluate LEREL’s performance. These datasets encompass diverse emotional categories, data dimensions, and signal characteristics, establishing a robust evaluation framework. The experimental results show that LEREL significantly outperforms the competing methods in key metrics such as recognition accuracy and F1-score, demonstrating its efficiency, robustness, and wide applicability.
3.3 Effects of Lipschitz Continuity Constraints
In our study, we applied Lipschitz continuity constraints across key components of the LEREL framework, including the EnhancedNASM component, attention mechanism, and normalization operations, to improve model stability and robustness. Experimental results (Figures 9 and 9) show that these constraints significantly enhanced training stability, particularly on low-accuracy datasets, comfirming their effectiveness for performance improvement.
Figure 8: EAV Sub2 with different Lipschitz constant
Figure 9: EAV Sub35 with different Lipschitz constant
3.4 Ablation Studies
The proposed LEREL framework employs an ensemble learning strategy comprising four independent sets of feature extractors, enhancers, and classifiers for EEG-based emotion recognition. Each set independently processes EEG signals to improve feature discriminability, while learnable weights optimize classifier integration (Table 4). Experimental results demonstrated that: 1) each classifier achieve relatively high accuracy. 2) the ensemble learning strategy significantly improves stability and robustness, and 3) ablation studies confirm its effectiveness in reducing misclassification rates.
Table 4: The results of ablation studies on the EAV dataset
Ablation model Average accuracy Average F1-score
Classifier 1 69.56 69.24
Classifier 2 68.23 67.90
Classifier 3 61.01 60.29
Classifier 4 62.10 61.72
Full model (average) 74.25 73.94
Full model (best) 76.43 76.22
3.5 Validation on the EAV Passive Dataset
The EAV dataset includes both active (recorded during conversations) and passive (collected during stimulus listening) signals. Passive signals present greater analytical challenges due to their higher noise levels, attenuated amplitude and weaker correlations with affective states. To evaluate model robustness, we specifically tested performance on passive signals. As shown in Figure 10, LEREL maintained a high classification accuracy under these suboptimal conditions, demonstrating exceptional signal processing stability.
Figure 10: Using passive data experiments to compare with other active data experiments on the EAV dataset
Figure 11: Partitioning with contextual information
Figure 12: Partitioning without contextual information
Figure 13: Real-time validation test of EAV
Figure 14: Real-time validation test of FACED
3.6 Real-time Validation Test
Real-time validation paradigms demand stricter performance criteria than traditional approaches. While conventional methods use complete contextual information through random sampling to optimize accuracy (Figure 12), real-time systems must operate solely on available signals without future context (Figure 12). This fundamental limitation necessitates superior generalization capacity and enhanced model robustness.
The proposed LEREL model demonstrates robust real-time emotion recognition, achieving consistently high classification accuracy on both EAV and FACED datasets (Figures 14 and 14). These results confirm its stronggeneralization capacity and practical applicability in complex real-world environments. Specifically, LEREL’s temporal dependency modeling enables effective EEG feature extraction and noise suppression, ensuring stable performance under challenging conditions.
4 Conclusion
This paper presents LEREL, a robust EEG-based emotion recognition framework that integrates ensemble learning with Lipschitz regularization. The framework combines: 1) Lipschitz constraints for noise-robust feature extraction, and 2) ensemble classifiers to optimize bias-variance trade-offs, ensuring reliable generalization. Experimental results demonstrate LEREL’s superior performance in both offline and real-time scenarios, particularly with challenging EEG signals. Comprehensive evaluations across multiple datasets and ablation studies validate each component’s contribution, showing consistent high accuracy while maintaining stability under noisy conditions. These results highlight LEREL’s practical potential for real-world affective computing. Future work will investigate computational optimization and multimodal integration to further enhance emotion recognition capabilities.